Could you please explain how to write this problem in an equation?
Bob is 4 years younger than 3 times Jerry's age. Together their ages add to 32. How old are Bob and Jerry?
My mother figured out that the answer is that Bob is 23 and Jerry is 9 but she did not know how to set it up in algebaic form.
I appreciate your help.

Answer by Paul(988)

Bob is 4 years younger than 3 times Jerry's age. Together their ages add to 32. How old are Bob and Jerry?
Jerry = x
Bob = 3x-4
x+(3x-4) = 32
4x-4+4=32+4
4x=36
4x/4=36/4
x=9
3x9-4 = 23
Therefore jerry is 9 years old and bob is 23 years old.